OpenStudy (anonymous):

Which of the following is the vertex of the function? y = x2 - 6x + 9 (3, 0) (-3, 36) (4, 1) (1, 4)

OpenStudy (jdoe0001):

you can always find the verrtex of one of those parabolas at \(\bf \left(-\cfrac{b}{2a}, c-\cfrac{b^2}{4a}\right)\)

OpenStudy (zzr0ck3r):

\[ax^2+bx+c\\\text{then the vertex is}\\\frac{-b}{2a}\]

OpenStudy (anonymous):

@mathstudent55

OpenStudy (zzr0ck3r):

what do you not understand? what is a,b,c in your equation?

OpenStudy (debbieg):

Actually, -b/(2a) is not the vertex, it's the x-coordinate of the vertex. Then find the y coordinate by evaluating the function at that x, e.g., f(-b/(2a))=y

OpenStudy (zzr0ck3r):

right, I was gonna get to that. baby steps:)

OpenStudy (zzr0ck3r):

I should have said the vertex is at* -b/2a

OpenStudy (debbieg):

lol... sorry @zzr0ck3r ... I always say, "If it seems like I'm being picky, it's because math is a very picky science" lol :)

OpenStudy (zzr0ck3r):

you are absolutely right